# LlamaLens: Specialized Multilingual LLM for Analyzing News and Social Media Content

### v1

@@ 4.2.2 Experimental Setup
The order of instructions in the training dataset can significantly impact model performance. For example, Pang et al. (2024) demonstrate that ordering instructions by complexity influences the effectiveness of tuned models. In light of this, we investigate how different orders of samples affect the performance of LlamaLens, employing four different data shuffling and ordering techniques to identify the optimal sequence. Although there are several possible ordering configurations, however, we limit the scope of this study to the effects of language and task order.
@@
For all models we train, we set a LoRA learning rate of 2e-4 alongside a linear learning rate schedule. Optimization was performed using AdamW Loshchilov and Hutter (2017), with a batch size of 16. All experiments were executed on four NVIDIA H100-80GB GPUs. In the first set of experiments, we trained four models quantized to 4-bit precision using QLoRA. Although the models store weights in 4-bit format, computations are performed in BFLOAT16 (bf16), with both the LoRA rank and $\alpha$ set to 16. Each model was trained on one of the dataset shuffling configurations. After identifying the optimal order of the training dataset—based on average model performance on our test sets, as shown in Figure 3—we used that dataset order to fine-tune our LlamaLens model in full precision (16-bit). Due to the scale of the model and training set size, we train the model for two epochs, increasing LoRA rank to 512, following the recommendations in Xin et al. (2024), which suggests that higher ranks yield improved performance for multitask learning. LoRA $\alpha$ was set equal to the rank.
@@ 4.3 Evaluation
4.4 Evaluation Process
For evaluation, we employed a zero-shot approach, in which we directly prompted the models (original and fine-tuned) to perform tasks from the testing sets. The employed natural instruction/prompt is the first generated instruction (Section 4.1) for each dataset. To ensure reproducibility, we set the top_p to 1 and the temperature to 0. For classification tasks, we restrict the models to generate a maximum of 20 new tokens. For summarization, we applied a stopping criterion based on encountering a period (‘.’) or the end-of-sequence (eos_token).
@@ 4.5 Post-processing
As models can generate text beyond that required in the instruction, a post-processing method was implemented to extract labels from the generated models’ responses. Initially, a regular expression was used to accurately identify and extract the labels. Several transformations were also applied, including converting all text to lowercase, removing special characters, and addressing code-switching by replacing non-Latin characters with their Latin equivalents. A similar post-processing approach was reported by Abdelali et al. (2024).
@@ 4.6 Evaluation Metrics
All models were evaluated using standard classification metrics including weighted-, macro-, micro- F1 and accuracy. We use ROUGE-2 to evaluate the summarization task. Specifically, we use the same metric reported in state-of-the-art (SOTA) per dataset.
Dataset Metric Llama Alpha. Full Task Lang. Dataset Metric Llama Alpha. Full Task Lang.
@@
CT24_T1 F1_Pos 0.404 0.538 0.583 0.893 0.288 claim-detection Mi-F1 0.545 0.895 0.891 0.884 0.898
ans-claim Ma-F1 0.335 0.427 0.528 0.454 0.473 Cyberbullying Acc 0.175 0.664 0.794 0.781 0.764
CT22Claim Acc 0.581 0.712 0.735 0.715 0.723 emotion Ma-F1 0.353 0.647 0.662 0.584 0.654
ArCyc_CB Acc 0.766 0.767 0.818 0.840 0.776 News_dataset Acc 0.654 0.502 0.712 0.787 0.614
Emotional-Tone W-F1 0.358 0.595 0.635 0.641 0.609 Politifact W-F1 0.121 0.210 0.241 0.252 0.262
NewsHeadline Acc 0.406 0.316 0.319 0.387 0.288 CNN_News_Articles Acc 0.644 0.897 0.919 0.904 0.911
Arafacts Mi-F1 0.210 0.376 0.263 0.466 0.362 News_Category Ma-F1 0.970 0.964 0.913 0.635 0.668
COVID19Factuality W-F1 0.492 0.794 0.733 0.595 0.780 SemEval23T3-ST1 Mi-F1 0.687 0.325 0.494 0.470 0.590
CT22Harmful F1_pos 0.507 0.539 0.565 0.473 0.539 xlsum R-2 0.074 0.088 0.126 0.116 0.101
annotated-hatetweets-4 W-F1 0.257 0.436 0.311 0.371 0.394 Offensive_Hateful Mi-F1 0.692 0.791 0.768 0.792 0.778
OSACT4SubtaskB Mi-F1 0.819 0.946 0.901 0.910 0.911 offensive_language Mi-F1 0.646 0.893 0.871 0.657 0.821
ASND Ma-F1 0.587 0.790 0.787 0.803 0.811 hate-offensive-speech Acc 0.602 0.874 0.901 0.909 0.903
SANADAkhbarona Acc 0.784 0.924 0.904 0.930 0.938 QProp Ma-F1 0.759 0.773 0.803 0.751 0.773
SANADAlArabiya Acc 0.893 0.975 0.973 0.973 0.973 News-Headlines Acc 0.668 0.959 0.961 0.953 0.960
SANADAlkhaleej Acc 0.865 0.929 0.920 0.916 0.929 NewsMTSC-dataset Ma-F1 0.628 0.640 0.669 0.685 0.613
UltimateDataset Ma-F1 0.376 0.742 0.594 0.647 0.673 CT24_T2 Ma-F1 0.535 0.464 0.440 0.554 0.379
NewsCredibility Acc 0.455 0.665 0.845 0.904 0.600 Hindi
xlsum R-2 0.034 0.058 0.058 0.061 0.063 fake-news Mi-F1 0.759 0.802 0.633 0.567 0.653
ArCyc_OFF Ma-F1 0.489 0.835 0.846 0.856 0.836 hate-speech-detection Mi-F1 0.750 0.910 0.898 0.903 0.879
OSACT4SubtaskA Ma-F1 0.782 0.876 0.852 0.863 0.849 Hindi-Hostility W-F1 0.469 0.666 0.564 0.664 0.526
ArPro Mi-F1 0.597 0.660 0.623 0.696 0.655 NLI_dataset W-F1 0.633 0.516 0.573 0.537 0.564
ArSarcasm-v2 F1_Pos 0.477 0.154 0.542 0.429 0.472 xlsum R-2 0.078 0.074 0.094 0.095 0.080
ar_reviews_100k F1_Pos 0.343 0.552 0.626 0.614 0.626 Offensive_Speech Mi-F1 0.621 0.692 0.701 0.733 0.763
ArSAS Acc 0.603 0.780 0.774 0.763 0.772 MC-Hinglish1.0 Acc 0.233 0.438 0.435 0.560 0.545
stance Ma-F1 0.608 0.634 0.774 0.775 0.853 Sentiment Analysis Acc 0.552 0.640 0.643 0.636 0.624
Mawqif-Arabic-Stance Ma-F1 0.764 0.774 0.819 0.845 0.846
ThatiAR F1_Pos 0.562 0.558 0.544 0.574 0.591
@@
Table 2: Quantized models performance across all datasets and dataset shuffling techniques. R-2: ROUGE-2. Acc: Accuracy. Mi-F1: Micro-averaged F1. Ma-F1: Macro-averaged F1. W-F1: weighted F1. F1_Pos: F1 of the positive class. Llama: Llama 3.1-8B-Instruct. Alpha.: Quantized model trained over dataset ordered alphabetically. Full: Quantized model trained over fully randomized dataset. Task: Quantized model trained over dataset shuffled by task. Lang.: Quantized model trained over dataset shuffled by language. Numbers in bold represent the best performance per dataset.
@@ 5 Results and Discussions
5.1 Data Shuffling Effect
In Figure 3, we present the results as averages computed across datasets and languages. Detailed results are provided in Table 2. Among the four different ordering strategies, shuffling by task outperformed other configurations. Both shuffling by language and alphabetic ordering showed comparable performance. Based on this experiment, we elect to train our LlamaLens model following the shuffling by task data ordering approach.
@@ 5.1 Data Shuffling Effect
5.2 Results on LlamaLens
In Table 1, we report the full results of our LlamaLens model across the different languages. Overall, LlamaLens significantly outperforms the Llama-instruct with an average improvement of 33%. Language wise LlamaLens outperforms Llama-instruct on average 32.5%, 32.2% and 29.5% for Arabic, English and Hindi, respectively. For Arabic, out of 26 tasks, LlamaLens only under-performs on cyberbullying, sarcasm and one of the hate-speech datasets. For English, out of 17 tasks LlamaLens outperforms on 14 tasks and lower performing tasks include sentiment, subjectivity and news categorization. For Hindi, in all 8 datasets LlamaLens outperforms Llama-instruct. In comparison to SOTA’s average performance of 0.75, LlamaLens average performance is 0.69. We should note here that since 18 out of the 52 datasets were not pre-split (Section 3.2), SOTA on these datasets is not directly comparable to our model, as the testing splits might be different. Computing average performance excluding these datasets reduces the gap, with a SOTA of 0.72 and LlamaLens performance of 0.67. In terms of number of datasets where LlamaLens improved over SOTA, we find that it outperforms SOTA in 16 test sets, and has a comparable performance (difference $\geq$ 0 and $<$-0.03) in 10 other testing sets. Dataset-specific improvements for English include tasks such as checkworthiness, emotion, news factuality, and propaganda. For Arabic, improvements have been observed in checkworthiness, emotion, and stance, while performance on news categorization is mixed, with improvements seen in 2 out of 5 datasets. For Hindi, significant improvements have been observed in hate speech, followed by NLI and offensive language.
5.3 Error Analysis of Llama-instruct Model
@@
We conducted an error analysis on the responses of the Llama-instruct model, to identify challenges in handling text of low-resource languages on certain classification tasks. Several common issues emerged across these models, particularly in tasks related to offensive language, and hate-speech detection, factuality, and news categorization. One recurring problem was the inability of the model to provide labels in numerous instances, often responding with phrases like “I cannot provide a label” or “Arabic text is not easily classifiable into categories without context or translation.” Another observed issue is language confusion in output. Although the models were instructed to output labels exclusively in English, they occasionally returned labels in Arabic or code-switched responses, which is inline with language confusion reported in relevant studies Marchisio et al. (2024), however, differently, we showcase the confusion can occur at the smallest unit of a single character. For instance, in some cases, the model generated outputs like “\<ف>actual” (referring to “factual” where Arabic character \<ف> is a transliteration of character “f”), and “\<س>포츠” (Korean for “sport” transliterated as “seupocheu” where character \<س> is a transliteration of character “s”) despite no instructions involving Korean language. This highlights a phonetic-level code-switching phenomenon. It also occurred on longer sequences such as the model responding with “\<سار>castic” instead of “Sarcastic,” where “\<سار>” is actually pronounced similarly to “Sar”. In contrast, our fine-tuned versions of the model do not display such issues, with the models adhering more consistently to the instructed output languages.This suggests that fine-tuning is critical for improving language-specific performance.

### v2

@@ 4.2.2 Experimental Setup
The order of instructions in the training dataset can significantly impact model performance. For example, Pang et al. (2024) demonstrate that ordering instructions by complexity influences the effectiveness of tuned models. In light of this, we investigate how different orders of samples affect the performance of LlamaLens, employing four different data shuffling and ordering techniques to identify the optimal sequence. Although multiple ordering configurations exist, we focus on the effects of language and task order.
@@
For all models we train, we set a LoRA learning rate of 2e-4. Optimization was performed using AdamW Loshchilov and Hutter (2017), with a batch size of 16. All experiments were executed on four NVIDIA H100-80GB GPUs. In the first set of experiments, we trained four models quantized to 4-bit precision using QLoRA. Although the models store weights in 4-bit format, computations are performed in BFLOAT16 (bf16), with both the LoRA rank and $\alpha$ set to 16. Each model was trained on one of the dataset shuffling configurations. After identifying the optimal order of the training dataset—based on average model performance on our test sets, as shown in Figure 3—we used that dataset order to fine-tune our LlamaLens model in full precision (16-bit). Due to the scale of the model and training set size, we train the model for two epochs, increasing LoRA rank to 128, following the recommendations in Xin et al. (2024), which suggests that higher ranks yield improved performance for multitask learning. LoRA $\alpha$ was set equal to the rank.
@@ 4.3 Evaluation
For the evaluation, we employed a zero-shot approach, in which we directly prompted the models to perform tasks from the testing sets. The employed natural instruction/prompt is the first generated instruction (Section 4.1) the temperature was set to 0 and top_p to 0.95. We manually checked a sample of instructions per task and found that they accurately expressed the intended task. [^6] Scripts are available at: https://github.com/firojalam/LlamaLens
@@ 4.4 Post-processing
As models can generate text beyond that is required in the instruction, a post-processing method was implemented to extract labels from the generated models’ responses. Initially, a regular expression was used to accurately identify and extract the labels. Several transformations were applied, including lowercasing, removing special characters, and handling code-switching by replacing non-Latin characters with Latin equivalents, similar to Abdelali et al. (2024); Dalvi et al. (2024).
@@ 4.5 Evaluation Metrics
All models were evaluated using standard classification metrics: weighted, macro, micro F1, and accuracy. Summarization was assessed with ROUGE-2, Specifically, we use the same metric reported in state-of-the-art (SOTA) per dataset.
Dataset Metric Model Performance Dataset Metric Model Performance
Base Alpha. Full Task Lang. Base Alpha. Full Task Lang.
@@
CT24_T1 F1_Pos 0.610 0.416 0.555 0.689 0.549 claim-detection Mi-F1 0.545 0.895 0.891 0.884 0.898
CT22Claim Acc 0.581 0.712 0.735 0.715 0.723 Cyberbullying Acc 0.175 0.664 0.794 0.781 0.764
ArCyc_CB Acc 0.766 0.767 0.818 0.840 0.776 emotion Ma-F1 0.353 0.647 0.662 0.584 0.654
Emotional-Tone W-F1 0.358 0.595 0.635 0.641 0.609 News_dataset Acc 0.654 0.502 0.712 0.787 0.614
NewsHeadline Acc 0.406 0.316 0.319 0.387 0.288 Politifact W-F1 0.121 0.210 0.241 0.252 0.262
Arafacts Mi-F1 0.210 0.376 0.263 0.466 0.362 CNN_News_Articles Acc 0.644 0.897 0.919 0.904 0.911
COVID19Factuality W-F1 0.492 0.794 0.733 0.595 0.780 News_Category Ma-F1 0.970 0.964 0.913 0.635 0.668
CT22Harmful F1_pos 0.507 0.539 0.565 0.473 0.539 SemEval23T3-ST1 Mi-F1 0.687 0.325 0.494 0.470 0.590
annotated-hatetweets-4 W-F1 0.257 0.436 0.311 0.371 0.394 xlsum R-2 0.074 0.088 0.126 0.116 0.101
OSACT4SubtaskB Mi-F1 0.819 0.946 0.901 0.910 0.911 Offensive_Hateful Mi-F1 0.692 0.791 0.768 0.792 0.778
ASND Ma-F1 0.587 0.790 0.787 0.803 0.811 offensive_language Mi-F1 0.646 0.893 0.871 0.657 0.821
SANADAkhbarona Acc 0.784 0.924 0.904 0.930 0.938 hate-offensive-speech Acc 0.602 0.874 0.901 0.909 0.903
SANADAlArabiya Acc 0.893 0.975 0.973 0.973 0.973 QProp Ma-F1 0.759 0.773 0.803 0.751 0.773
SANADAlkhaleej Acc 0.865 0.929 0.920 0.916 0.929 News-Headlines Acc 0.668 0.959 0.961 0.953 0.960
UltimateDataset Ma-F1 0.376 0.742 0.594 0.647 0.673 NewsMTSC-dataset Ma-F1 0.628 0.640 0.669 0.685 0.613
NewsCredibility Acc 0.455 0.665 0.845 0.904 0.600 CT24_T2 Ma-F1 0.535 0.464 0.440 0.554 0.379
ArCyc_OFF Ma-F1 0.489 0.835 0.846 0.856 0.836 Average 0.528 0.629 0.673 0.662 0.620
xlsum R-2 0.034 0.058 0.058 0.061 0.063 Hindi
OSACT4SubtaskA Ma-F1 0.782 0.876 0.852 0.863 0.849 fake-news Mi-F1 0.759 0.802 0.633 0.567 0.653
ArPro Mi-F1 0.597 0.660 0.623 0.696 0.655 hate-speech-detection Mi-F1 0.750 0.910 0.898 0.903 0.879
ArSarcasm-v2 F1_Pos 0.477 0.154 0.542 0.429 0.472 Hindi-Hostility W-F1 0.469 0.666 0.564 0.664 0.526
ar_reviews_100k F1_Pos 0.681 0.552 0.626 0.614 0.626 NLI_dataset W-F1 0.633 0.516 0.573 0.537 0.564
ArSAS Acc 0.603 0.780 0.774 0.763 0.772 xlsum R-2 0.078 0.074 0.094 0.095 0.080
stance Ma-F1 0.608 0.634 0.774 0.775 0.853 Offensive_Speech Mi-F1 0.621 0.692 0.701 0.733 0.763
Mawqif-Arabic-Stance Ma-F1 0.764 0.774 0.819 0.845 0.846 MC-Hinglish1.0 Acc 0.233 0.640 0.643 0.636 0.545
ThatiAR F1_Pos 0.562 0.558 0.544 0.574 0.591 Sentiment Analysis Acc 0.552 0.627 0.650 0.658 0.624
Average 0.540 0.636 0.653 0.668 0.659 Average 0.4777 0.589 0.589 0.604 0.569
@@
Table 2: Performance of models trained with QLoRA across all datasets and dataset shuffling techniques. R-2: ROUGE-2. Acc: Accuracy. Mi-F1: Micro-averaged F1. Ma-F1: Macro-averaged F1. W-F1: weighted F1. F1_Pos: F1 of the positive class. Base: Llama 3.1-8B-Instruct. Alpha: Model trained with QLoRA on a dataset ordered alphabetically. Full: Model trained with QLoRA on a fully randomized dataset. Task: Model trained with QLoRA on a dataset shuffled by task. Lang.: Model trained with QLoRA on a dataset shuffled by language. Numbers in bold indicate the best performance per dataset.
@@ 5 Results and Discussions
To contextualize the performance of LlamaLens, we compare it against both the SOTA results and the Llama-instruct baseline. SOTA serves as a theoretical upper bound, representing the best-reported results on each dataset, while the Llama-instruct model acts as a lower bound, reflecting a general-purpose instruction-tuned LLM without task-specific adaptation. Our evaluation aims to assess how well LlamaLens bridges the gap between these two extremes, demonstrating its ability to generalize while leveraging task-specific fine-tuning.
5.1 LlamaLens Performance
In Table 1, we report the full results of our LlamaLens model across the different languages. Overall, LlamaLens significantly outperforms the Llama-instruct with an average improvement of 62%. Language wise LlamaLens outperforms Llama-instruct on average 59%, 71% and 56% for Arabic, English and Hindi, respectively. For Arabic, LlamaLens demonstrates strong performance across 27 tasks, underperforming only in Checkworthiness and one Subjectivity dataset. In English, LlamaLens outperforms in 15 out of 17 tasks, with News Categorization being the primary area of lower performance. For Hindi, across 8 datasets, LlamaLens surpasses Llama-Instruct in 7, with Natural Language Inference being the only dataset where it under-performs compared to Llama-Instruct. Comparable Performance with SOTA. In comparison to SOTA’s average performance of 0.75, LlamaLens achieves an average performance of 0.727. We should note that since 18 out of the 52 datasets were not pre-split (Section 3.2), SOTA on these datasets is not directly comparable to our model, as the testing splits might differ. For some datasets, issues such as duplicate entries, missing input text, or the absence of development or test splits (e.g., CNN_News_Articles for English) required us to clean the dataset and create new splits. As a result, a direct apple-to-apple comparison may not always be possible. However, the reported SOTA scores serve as a close approximation for meaningful evaluation. Computing the average performance excluding these datasets reduces the gap, with a SOTA of 0.716 and LlamaLens performance of 0.693. In terms of number of datasets where LlamaLens improved over SOTA, we find that it outperforms SOTA in 23 test sets, and has a comparable performance (difference between -0.04 to 0) in 8 other testing sets. Performance Gains Across Datasets. Dataset-specific improvements in English include gains in Factuality, with Propaganda and Checkworthiness showing significant advancements. Summarization improved across both English and Hindi, along with News Genre Categorization (two datasets in English and four in Arabic). In Hindi, Offensive Language demonstrated notable improvements, while Hate Speech and Cyberbullying also exhibited gains, with the latter performing better in Arabic as well. Additionally, Arabic showed stronger performance in News Credibility, Emotion, Stance, Attentionworthiness, and Claim Detection.
5.2 Native vs. English Instructions
When comparing the two versions of LlamaLens (trained with English vs. Native instructions), the performance difference between them is minimal. The results are closely aligned, with the Native model [^7] The model was trained using language-specific instructions tailored to the dataset. achieving an average score of 0.723, compared to 0.727 for the English-instructed model. A closer look reveals that the English-instructed model outperformed the Native model in 28 datasets, while the Native model led in 24. Overall, 49 of 52 datasets showed comparable performance, with differences between -0.05 and 0.05. The most notable difference was in News_Category_Dataset, where the English-instructed model outperformed the Native model by 0.304.
5.3 Impact of Data Shuffling
Figure 3 shows the averaged results across datasets and languages, with details in Table 2. Shuffling by task achieved the highest performance, while shuffling by language and alphabetic ordering performed similarly but did not match the effectiveness of the task-based approach. To determine statistical significance, we performed a Wilcoxon signed-rank test comparing the best (shuffled by task) and worst (alphabetical) configurations. The test yielded a p-value of 0.025, below the $\alpha=0.05$ threshold, confirming that improvements from task-based shuffling are not random. Based on this, we adopt shuffling by task for training LlamaLens to enhance performance across diverse datasets.
@@ 5.3 Impact of Data Shuffling
5.4 Task-wise Results
We computed task-wise performance differences, as detailed in Table 3. Fine-tuned models demonstrated significant improvements, particularly in English Cyberbullying tasks, where LlamaLens achieved a performance gain of more than 61% compared to other languages. Across all three languages, LlamaLens consistently outperformed baselines in sentiment analysis, summarization, factuality, hate speech, and offensive language detection. Notably, when evaluated across 37 combined datasets, LlamaLens underperformed in only four cases: natural language inference, subjectivity (Arabic), checkworthiness (Arabic), and news genre categorization (specifically the SemEval23T3-ST1 dataset in English).
Task Lang #DS Base L-Lens (Eng.) $\Delta$
Summarization Arabic 1 0.034 0.129 0.095
English 1 0.074 0.182 0.108
Hindi 1 0.078 0.171 0.093
News Cat. Arabic 5 0.701 0.942 0.241
English 3 0.767 0.678 -0.088
Emotion Arabic 2 0.382 0.592 0.211
English 1 0.353 0.803 0.450
Sarcasm Arabic 1 0.477 0.520 0.043
English 1 0.668 0.936 0.268
Sentiment Arabic 2 0.642 0.792 0.150
English 1 0.628 0.751 0.123
Hindi 1 0.552 0.647 0.095
Stance Arabic 2 0.686 0.890 0.203
News Credibility Arabic 1 0.455 0.935 0.480
Attentionworthy Arabic 1 0.158 0.425 0.267
Checkworthiness Arabic 1 0.610 0.425 -0.185
English 1 0.404 0.942 0.539
Claim Arabic 1 0.581 0.734 0.153
English 1 0.545 0.864 0.319
Factuality Arabic 2 0.351 0.785 0.434
English 2 0.387 0.643 0.256
Hindi 1 0.759 0.994 0.235
Propaganda Arabic 1 0.597 0.747 0.150
English 1 0.759 0.963 0.204
Cyberbullying Arabic 1 0.766 0.870 0.104
English 1 0.175 0.836 0.661
Hindi 1 0.233 0.625 0.392
Harmfulness Arabic 1 0.507 0.523 0.016
Hate Speech Arabic 2 0.538 0.740 0.203
English 1 0.602 0.931 0.329
Hindi 2 0.609 0.858 0.249
Offensive Arabic 2 0.636 0.887 0.251
English 3 0.647 0.881 0.235
Hindi 1 0.621 0.862 0.241
Subjectivity English 1 0.535 0.642 0.107
Arabic 1 0.562 0.441 -0.121
NLI Hindi 1 0.633 0.568 -0.065
Table 3: Task-Based Evaluation Across Arabic, English, and Hindi. Lang: Languages. #DS: Number of Datasets. Base: Llama3.1-Instruct. L-Lens: LlamaLens. Eng: English. News Cat: News Genre Categorization. NLI: Natural Language Inference. The $\Delta$ Column represents the difference between two models.
5.5 Capability-based Results
Table 9 presents LlamaLens’s performance across key task categories for Arabic, while Table 10 covers English and Hindi. These categories include Natural Language Understanding (NLU), Information Extraction and Text Classification, Emotion and Sentiment Analysis, Fact-Checking and Verification, and Hate Speech and Offensive Content, as illustrated in Figure 1. The tables report the SOTA scores, the Llama-instruct baseline, and LlamaLens using native instructions. LlamaLens achieves notable gains over Llama-instruct, particularly in Arabic and English, with English tasks consistently performing better due to the availability of abundant datasets and established benchmarks. Despite being a medium-resource language, Arabic shows significant improvements, closing the performance gap with English across multiple tasks. For Hindi, while overall scores remain lower, LlamaLens demonstrates clear advancements, particularly in emotion and sentiment analysis and hate speech and offensive content, highlighting its potential in underrepresented languages. Hate speech and offensive Content emerges as LlamaLens’s strongest task across all three languages, with the highest improvement recorded in English (+0.411), demonstrating its capability in handling complex linguistic challenges.
5.6 Error Analysis
@@
We analyzed Llama-Instruct’s responses to identify challenges across different tasks, such as handling offensive language, hate speech, factuality, and news categorization. Model Hesitation and Contextual Dependence. One recurring problem was the inability of the model to provide labels in numerous instances, often responding with phrases like “I cannot provide a label” or “Arabic text is not easily classifiable into categories without context or translation.” Such responses occur when an LLM lacks context to classify text confidently or is designed to avoid labeling sensitive topics (e.g., political, religious, or controversial). These hesitations stem from built-in safeguards to prevent incorrect classifications. Language Confusion. Another observed issue is language confusion in output. Although the models were instructed to output labels exclusively in English, they occasionally returned labels in Arabic or code-switched responses, which is inline with language confusion reported in relevant studies Marchisio et al. (2024), however, differently, we showcase the confusion can occur at the smallest unit of a single character. For instance, in some cases, the model generated outputs like “\<ف>actual” (referring to “factual” where Arabic character \<ف> is a transliteration of character “f”), and “\<س>포츠” (Korean for “sport” transliterated as “seupocheu” where character \<س> is a transliteration of character “s”) despite no instructions involving Korean language. This highlights a phonetic-level code-switching phenomenon. It also occurred on longer sequences such as the model responding with “\<سار>castic” instead of “Sarcastic,” where “\<سار>” is actually pronounced similarly to “Sar”. In contrast, our fine-tuned versions of the model do not display such issues. This suggests that fine-tuning is critical for improving language-specific performance.